OpenStudy (anonymous):

if h(x)=x^2-1, and h(2a)=35, what is the value of a?

OpenStudy (mrnood):

h(2a) = (2a)^2-1 = 35 solve for a

OpenStudy (anonymous):

@MrNood i got 9 and the answer is a= {3,-3}

OpenStudy (mrnood):

4a^2= 36 a^2=9 a= +/- 3 I'm not sure about your notation (do you mean the set of {3,-3}?)

OpenStudy (anonymous):

yeah the set of

OpenStudy (mrnood):

good - well done. Do you understand how I went from h(x) to h(2a)?

OpenStudy (anonymous):

oh yeah i got it i just forgot the ^2 part thanks!